OpenStudy (anonymous):

MEDAL AND FAN The following two ratios are equivalent. Find the missing value, 4/6 = x/21

OpenStudy (er.mohd.amir):

4*21=x*6 6x=84 x=84/6 x=14

OpenStudy (anonymous):

@Er.Mohd.AMIR Can you explain how you got there? What you are doing? 1) That's the point of this website. 2) I can't follow what you just did.

OpenStudy (anonymous):

I see how he got it.